OpenStudy (fanduekisses):

I'm having trouble with this equation: (x-5)/2 + (1/2)=2x-(x-3)/8

OpenStudy (fanduekisses):

\[\frac{ x-5 }{ 2} + \frac{ 1 }{ 2 }= 2x-\frac{ x-3 }{ 8}\]

OpenStudy (anonymous):

Remove all denominators by multiplying all the terms by the LCD, which is 8

OpenStudy (fanduekisses):

ok, so it will be x-5/4?

OpenStudy (anonymous):

It will be: 4(x-5) + 4 = 16 x - (x - 3)

OpenStudy (anonymous):

you can continue from there

OpenStudy (fanduekisses):

ohhh so i multiply all the terms by 8, but not the variables

OpenStudy (fanduekisses):

I see :) thanks so much!

OpenStudy (fanduekisses):

Then I'll just distribute and combine like terms right?

OpenStudy (anonymous):

Yes.

OpenStudy (fanduekisses):

thanks

OpenStudy (fanduekisses):

I got x=\[-\frac{ 19 }{ 11 }\] is this correct?